### v1

@@ 4.6 Power Dependent Mechanical Linewidth
An increase in the intrinsic linewidth from to is observed following the cladding of the OMC in alumina. Both of these values are two to three orders of magnitude narrower than the corresponding frequency feature observed in the spin decay spectrum measurement in Fig. 3. We attribute this difference to excess mechanical dissipation introduced by the gas condensation procedure required to achieve the optical resonance condition between the TE optical mode and the SiV center’s ZPL.
@@ 4.8 Temperature-Dependent Spin Decay
Fig. 13 shows the temperature-dependent spin relaxation data measured by applying a controlled amount of heat to the mixing plate of the dilution fridge. As the mixing plate temperature is raised, the spin qubit is found to decay faster with a steady state () population value tending towards 0.5 which corresponds to a completely mixed spin state. This behavior is expected at higher temperatures relative to the spin-transition frequency. The SiV temperature estimate for each measurement can be extracted from the long-time population saturation value and is shown inset to Fig. 13. In the temperature range probed in these measurements, the SiV temperature estimate from this method is in good agreement with the temperature reading on the mixing plate thermometer.
@@ 4.9 Determination of Extrinsic Optical Cavity Loss Rate
It was observed that the quality factor and depth of resonance dip changed after the device was cleaned and cladded in ALD alumina. We suspect but do not know that the cleaning was the step responsible for the slight increase. Fits to both spectra assuming an undercoupled cavity yields GHz. Fits assuming an overcoupled system do not yield consistent results for leading us to believe that our devices are undercoupled with GHz and GHz for the ALD alumina-cladded case. This is the best measurement of that we can do in the absence of cavity optical phase response measurements, which are difficult due to the relatively large optical linewidths of our device relative to the available bandwidth of visible-wavelength electro-optic modulators and photodetectors.
@@ 4.11 SEM Analysis of fabricated device
The goal of the OMC cavity design in this work is to achieve an optical/mechanical cavity mode pair which simultaneously enables optical readout of the SiV center’s spin properties and optomechanical readout of the mechanical mode properties. This requires high values of and low values of . There has been a great deal of previous work on designing mode pairs which satisfy these criteria.[60] We choose the fundamental TE-like optical mode, mechanical ‘breathing mode’ pair, which features a near optimal photoelastic overlap integral between the optical field component with the strain field component, which enables large optomechanical coupling. The TE-like optical mode also supports efficient spin-photon interactions and the mechanical breathing mode supports efficient spin-phonon interactions for an SiV oriented relative to the beam.
The cavity region of the fabricated device is shown in Fig 16. A unit cell design akin to the conventional elliptical hole shape is used (where the hole is defined by its major/minor axis lengths) except an additional parameter is introduced which determines the radius of curvature near the hole edge.
@@ 4.12 Simulation of the optomechanical cavity
The optical TE-like and the mechanical breathing modes were simulated using the geometric parameters obtained from SEM images. Fig 1(a) and Fig 1(b) show the simulated profiles of these modes. The masked implantation process embeds SiV centers within a 100 nm x 100 nm area (green square shown in Fig 17a and Fig 17b). The TE-like mode has a simulated wavelength of 736 nm and a quality factor of 3.2 104. The breathing mode has a simulated frequency of 12.29 GHz and a quality factor of 2.4 105. While the measured optical Q-factor is similar to the simulated quality factor, the measured mechanical Q-factor is much lower than the simulated value.

### v2

@@ 5.6 Power Dependent Mechanical Linewidth
An increase in intrinsic linewidth from to is observed following the cladding of the OMC in alumina. Both of these values are two to three orders of magnitude narrower than the corresponding frequency feature observed in the spin decay spectrum measurement in Fig. 3. We attribute this difference to excess mechanical dissipation introduced by the gas condensation procedure required to achieve the optical resonance condition between the TE-like optical mode and the SiV center’s ZPL.
@@ 5.8 Temperature-Dependent Spin Decay
Fig. 13 shows the temperature-dependent spin relaxation data measured by applying a controlled amount of heat to the mixing plate of the dilution fridge. As the mixing plate temperature is raised, the spin qubit is found to decay faster with a steady state () population value tending towards 0.5 which corresponds to a completely mixed spin state. This behavior is expected at higher temperatures relative to the spin transition frequency. The SiV temperature estimate for each measurement can be extracted from the long-time population saturation value and is shown inset to Fig. 13. In the temperature range probed in these measurements, the SiV temperature estimate from this method is in good agreement with the temperature reading on the mixing plate thermometer.
@@ 5.9 Determination of Extrinsic Optical Cavity Loss Rate
It was observed that the quality factor and depth of resonance dip changed after the device was cleaned and cladded in ALD alumina. We suspect, but do not know, that the cleaning was the step responsible for the slight increase. Fits to both spectra assuming an undercoupled cavity yields GHz. Fits assuming an overcoupled system do not yield consistent results for leading us to believe that our devices are undercoupled with GHz and GHz for the ALD alumina-cladded case. This is the best measurement of that we can make in the absence of cavity optical phase response measurements, which are difficult due to the relatively large optical linewidths of our device relative to the available bandwidth of visible-wavelength electro-optic modulators and photodetectors.
@@ 5.11 SEM Analysis of fabricated device
The goal of the OMC cavity design in this work is to achieve an optical/mechanical cavity mode pair which simultaneously enables optical readout of the SiV center’s spin properties and optomechanical readout of the mechanical mode properties. This requires high values of and low values of . There has been a great deal of previous work on designing mode pairs which satisfy these criteria.[63] We choose the fundamental TE-like optical mode, mechanical ‘breathing mode’ pair, which features a near-optimal photoelastic overlap integral between the optical field component with the strain field component, which enables large optomechanical coupling. The TE-like optical mode also supports efficient spin-photon interactions and the mechanical breathing mode supports efficient spin-phonon interactions for an SiV oriented relative to the beam.
The cavity region of the fabricated device is shown in Fig. 16. A unit cell design akin to the conventional elliptical hole shape is used (where the hole is defined by its major/minor axis lengths) except an additional parameter is introduced, which determines the radius of curvature near the hole edge.
@@ 5.12 Simulation of the optomechanical cavity
The optical TE-like and mechanical breathing modes were simulated using the geometric parameters obtained from SEM images. Fig 1(a) and Fig 1(b) show the simulated profiles of these modes. The masked implantation process embeds SiV centers within a 100 nm x 100 nm area (green square shown in Fig 17a and Fig 17b). The TE-like mode has a simulated wavelength of 736 nm and a quality factor of 3.2 104. The breathing mode has a simulated frequency of 12.29 GHz and a quality factor of 2.4 105. While the measured optical Q-factor is similar to the simulated quality factor, the measured mechanical Q-factor is much lower than the simulated value.